A normal population has a mean of 0.1 and standard deviation 2.1. Find the
probability that the mean of a sample of size 900 will be negative
1
Expert's answer
2019-03-18T13:35:31-0400

Given


"\mu=0.1, \sigma=2.1, n=900"


The standard normal variate is


"z=(xmean - \mu)/(\sigma/sqrt(n))""z=(xmean - 0.1)\/(2.1\/sqrt(900))=(xmean-0.1)\/0.07""xmean=0.1+0.07z"


The required probability, that the sample mean is negative is given by


"P(xmean<0)=P(0.1+0.07z<0)=P(z<-1\/0.7)=P(z<-1.42857)=0.076564"